OpenStudy (timaashorty):

I need help finding the variance and the standard deviation of some values. Can someone please take me through it?

OpenStudy (timaashorty):

Round to the nearest tenth. Find the mean, variance, & standard deviation of these values:

OpenStudy (timaashorty):

1,3,11,11,11,11

OpenStudy (timaashorty):

@doc.brown can you help?

OpenStudy (doc.brown):

Mean= average. What's the average?

OpenStudy (timaashorty):

1+3 + 11+11+11+11 = 48/ 6 = 8 Mean = 8

OpenStudy (timaashorty):

I need help in the variance and standard deviation

OpenStudy (doc.brown):

Variance is the average of the squares of the difference.

OpenStudy (timaashorty):

11-8 ^2 like that?

OpenStudy (doc.brown):

So, find out how far each value is from the mean. 1 is 8-1 so 7 away from the mean 3 is 8-3 = 5 etc

OpenStudy (timaashorty):

11 - 8 or 8 - 11 for the 11's?

OpenStudy (doc.brown):

11-8

OpenStudy (timaashorty):

11-8=3 11-8=3 11-8 =3 11- 8 = 3 8-1 = 7 8-3 =5

OpenStudy (timaashorty):

what's next?

OpenStudy (doc.brown):

Square them all, then find the average.

OpenStudy (timaashorty):

3^2 = 9 3^2 = 9 3^2 = 9 3^2 = 9 7^2 = 49 5^2 = 25

OpenStudy (timaashorty):

9*4 + 25+49 = 110/ 6 = 18.3

OpenStudy (timaashorty):

thats the variance correct?

OpenStudy (doc.brown):

That's the variance. 18.3 I can go over why it gets squared if you'd like. But for now take the square root of that for the Standard Deviation.

OpenStudy (timaashorty):

sqrt 18.3 = 4.278

OpenStudy (doc.brown):

So the Standard Deviation is 4.3 Which makes sense, your data values are basically 8.0\(\pm\)4.3

OpenStudy (timaashorty):

which equals 4.3

OpenStudy (timaashorty):

oh my thank you for taking your time to help me and everything!

OpenStudy (doc.brown):

No prob

OpenStudy (timaashorty):

Just to make sure mean: 8 Variance: 18.3 Standard deviation: 4.3

OpenStudy (timaashorty):

correct?

OpenStudy (doc.brown):

Round to the nearest tenth. 8.0

OpenStudy (timaashorty):

for the mean?

OpenStudy (doc.brown):

Yeah, it's a good habit and will save you marks later in life. Mean: 8.0 Variance: 18.3 Standard deviation: 4.3